# Numerical Ergodicity of Stochastic Allen–Cahn Equation driven by Multiplicative White Noise

### v1

@@ 4. Numerical Experiments
In this section, we perform a numerical experiment to verify our theoretical result, Theorem 3.2 in Section 4.
The numerical test is given to the stochastic Allen–Cahn equation (2.4) in $(0,1)=(0,1)$ with $\epsilon=0.5$ and $g(x)=2+\sin x^{2}$. By Theorem 3.2, the DIEG scheme (die-g) applied to Eq. (2.4) is uniquely ergodic for any $\tau\in(0,1)$ (fulfilling the conditon $(\epsilon^{-2}-\lambda_{1})\tau<1$). We take $\tau=0.05$ and $N=10$ (the dimension of the spectral Galerkin approximate space), choose three initial data $X_{0}(\xi)=\sin\pi\xi,\sum_{k=1}^{10}\sin k\pi\xi,-\sum_{k=1}^{10}\sin k\pi\xi$, $\xi\in(0,1)$, respectively, and approximate the expectation by taking averaged value over $1,000$ paths to implement the numerical experiments. In addition, we simulate the time averages $\frac{1}{2,000}\sum_{n=1}^{2,000}\mathbb{E}[\phi(X_{n}^{N})]$ (up to $n=2,000$ corresponding to $t=100$) by
@@
Figure 1. Time averages of (die-g) for Eq. (2.4)
From Figure 1, the time averages of the DIEG scheme (die-g) with different initial data converge to the same ergodic limit, which coincides with the theoretical result in Theorem 3.2. This experiment also indicates that the original 1D stochastic Allen–Cahn equation (2.4), with bounded, invertible, and continuous diffusion coefficient, driven by multiplicative white noise, is uniquely ergodic, which will be investigated in a separate paper.

### v2

@@ 4. Numerical Experiments
In this section, we perform some numerical experiments to verify our theoretical result, Theorem 3.2 in Section 3.
The numerical test is given to the stochastic Allen–Cahn equation (2.4) in $\mathscr{O}=(0,1)$ with $\epsilon=0.5$ and $g(x)=2+\sin x^{2}$. By Theorem 3.2, the DIEG scheme (DIEG) applied to Eq. (2.4) is uniquely ergodic for any $\tau\in(0,1)$ (fulfilling the conditon $(\epsilon^{-2}-\lambda_{1})\tau<1$). We take $\tau=0.05$ and $N=10$ (the dimension of the spectral Galerkin approximate space), choose three initial data $X_{0}(\xi)=\sin\pi\xi,\sum_{k=1}^{10}\sin k\pi\xi,-\sum_{k=1}^{10}\sin k\pi\xi$, $\xi\in(0,1)$, respectively, and approximate the expectation by taking averaged value over $1,000$ paths to implement the numerical experiments. In addition, we simulate the time averages $\frac{1}{2,000}\sum_{n=1}^{2,000}\mathbb{E}[\phi(X_{n}^{N})]$ (up to $n=2,000$ corresponding to $t=100$) by
@@
Figure 1. Time averages of (DIEG) for Eq. (2.4)
From Figure 1, the time averages of the DIEG scheme (DIEG) with different initial data converge to the same ergodic limit, which coincides with the theoretical result in Theorem 3.2. This experiment also indicates that the original 1D stochastic Allen–Cahn equation (2.4), with bounded, invertible, and continuous diffusion coefficient, driven by multiplicative white noise, is uniquely ergodic, which will be investigated in a separate paper.